1. The Anderson Palmer Hospital would like to reduce the cost of its inventory by calculating the optimal order size for hypodermic needles. On average, the hospital uses 200 needles per month, the cost of placing an order is $12, and holding cost is estimated to be $1.20 per needle per year. Using the EOQ formula, the optimal order size is:

about 60 needles.

about 140 needles.

about 220 needles.

about 310 needles.

2. The Anderson Palmer Hospital would like to reduce the cost of its inventory by calculating the optimal order size for hypodermic needles. On average, the hospital uses 200 needles per month. The purchasing staff has calculated that the EOQ for needles is 160 needles. If they proceed to purchase in EOQ quantities, how many orders per year will be placed?

About 15 orders per year

About 18 orders per year

About 20 orders per year

About 30 orders per year
Answer #1

Answer:

1)

Annual Demand D = 200 * 12 = 2400 needles

Order cost S = $12

Inventory holding cost H = $1.20

EOQ = sqrt [(2 * D * S) / H]

EOQ = sqrt [(2 * 2400 * 12) / 1.2]

   =sqrt (48000)

   =219.08
  
EOQ = 219.08 needles

=About 220 needles (Option C)

2)

EOQ = 160 needles

No of needles used by the hospital per month = 200

No of needles needed by the hospital in a year = 200*12 = 2400

No of orders = No of needles required annually / EOQ

= 2400 / 160

= 15

Since orders can’t be in a fraction

About 15 orders per year